• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 08:02:31 on March30,2005

    Entry number 142778

    keyword=Owl shift summary

    Owl Shift Summary
    ----------------------
    Production Runs on Tilted Carbon
    
    Run 3057 - 00:12-00:28  64k events
    
    00:28 MCC having problems with RF cavities, so expect beam to be down for 5-20 
    minutes.
    
    00:32 Beam back, so start DAQ again.
    
    Run 3058 - 00:32-01:30  241k events
    
    Run 3059 - 01:31-02:17  201k events
    
    Run 3060 - 02:18-03:03  204k events
    
    Run 3061 - 03:03-04:07  280k events
    
    DAQ problem - TS2 reboot, Also restarted CODA.  Run 3062 has no data in it for 
    this reason.
    
    Run 3063: 04:11-05:15  277k events *end of run failed (ER1 not reporting 
    status).  Needed to restart CODA and reboot all trigger supervisors.
    
    Run 3064-3066 - checking DAQ - no data.
    
    Run 3067 - 05:20-06:16  249k events
    
    Run 3068 - 06:17-07:23  270k events
    
    Run 3069 - 07:23- (still going)